OpenStudy (anonymous):

r(x)=3x+5/x^1+1, s(x)=x^2+3/3x-1 write the expression (2r-3x)(x) as a simplified ratio with the numerator and denominator each written as a sum of terms of the form cx^m and c>0, for the term in the highest power in the numerator (2r-3s)(x)=f(x)/g(x) where f(x)=? and g(x)=?

OpenStudy (anonymous):

is that really \[x^1+1\] in the denominator of the first one?

OpenStudy (anonymous):

let me doucle check

OpenStudy (anonymous):

x^2+1

OpenStudy (anonymous):

\[2\frac{3x+5}{x^2+1}=\frac{6x+10}{x^2+1}\]

OpenStudy (anonymous):

\[3s(x)=3\frac{x^2+3}{3x-1}=\frac{3x^2+9}{3x-1}\]

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

subtract and get \[\frac{-3 x^4+6 x^2+24 x-19}{(3 x-1) (x^2+1)}\]

OpenStudy (anonymous):

the way you answered, I'm slightly confused. could you put the answer in f(x) and g(x) form?

OpenStudy (anonymous):

assuming my algebra is correct, \[f(x)=\text{numerator}\] \[g(x)=\text{denominator}\]

OpenStudy (anonymous):

\[f(x)=-3 x^4+6 x^2+24 x-19\] in other words